# Domain Transfer Through Image-to-Image Translation for Uncertainty-Aware Prostate Cancer Classification

### v1

@@ 6.4 Filtering during deployment
So far, we explored the effect of the data filtering strategy during training to improve the model robustness. It is also possible to apply filtering on the test set, i.e. when deploying the model to real clinical routines. This is equivalent to refraining from making decisions on the test samples that are identified as highly uncertain. We use the pre-trained MpMRI and M.S. MpMRI models, each with 0% and 20% training filtering rate as final models and evaluate their performance on the test set. Figure 6 shows the experiment on filtering 0% to 40% of the test data when deploying the pre-trained models with 20% filtering during training. The performance of the other two models (0% filtering) can be found in Appendix B. We observe that the model improves its performance when filtering out highly uncertain patients from the test set, and eventually classified all patients correctly, as shown in Figure 5(a).
This approach has practical applications under real clinical settings, as it can help radiologists save much time by focusing on patients that have been filtered out (with high uncertainty value) during the diagnostic process, rather than well-classified patients.
((a))
((b))
Figure 6: Performance of filtering 0% to 40% of test data on the selected models. 5(a) represents the test performance of “MpMRI” model with 20% filtration on the training set, and 5(b) represents the test performance of “M.S. MpMRI” model with 20% filtration on the training set.
@@ 7 Conclusion
In this work, we presented a novel approach for unpaired image-to-image translation of prostate mp-MRI and developed a robust deep-learning model for classifying clinically significant PCa using evidential focal loss. We demonstrated the effectiveness of our method on our local dataset, reinforced by a publicly available one, and showed that uncertainty-aware filtering during both training and deployment can significantly improve the PCa classification performance. Out method has the potential to assist with and expedite the diagnostic process by suggesting highly uncertain patients where clinicians can focus on precise diagnosis and fast track those with high prediction certainty.
While our approach has shown promising results, there are still opportunities for improvement. One potential area for future work is to consider the spatial dependency between slices in volumetric MRI. Currently, our domain transfer framework only accepts 2D images as input and output, and we reshape the volumetric MRI into several 2D slices. However, explicitly splitting 3D images into 2D slices may eliminate the spatial dependency within each MRI data and affect the classification results. Therefore, one possible solution is to translate the 3D MRI as a whole from 3.0T to 1.5T instead of translating a single slice at a time.
[…]
@@ Acknowledgments
Ethical Standards
All procedures performed in studies involving human participants were in accordance with the ethical standards of the institutional and/or national research committee and with the 1964 Helsinki Declaration and its later amendments or comparable ethical standards.
Conflicts of Interest
We declare we don’t have conflicts of interest.
A Original Results
In this section, we provide all original results based on image patches (Patch-based results) of all experiments we performed in the paper. See Table 3 for the main results.
Data Rate Method Acc. Sen. Spec. AUC
SDNet(baseline) T2 0% N/A 77.8 27.9 90.7 74.8$\pm$4.1
T2-only T2 0% N/A 65.6 75.7 63.0 77.9$\pm$3.9
MpMRI T2+ADC 0% N/A 75.1 69.3 76.7 79.9$\pm$3.8
Vol. MpMRI T2+ADC 0% N/A 64.9 62.9 65.4 63.4$\pm$5.5
M.S. MpMRI T2+ADC 0% N/A 70.4 74.3 69.4 80.1$\pm$3.6
MpMRI+co-teaching T2+ADC 0% N/A 78.4 72.1 80.0 80.7$\pm$3.9
MpMRI T2+ADC 10% patch 77.5 80.7 76.7 82.4$\pm$3.4
M.S.MpMRI T2+ADC 10% patch 79.7 70.7 82.0 81.1$\pm$3.8
MpMRI T2+ADC 20% patch 83.8 80.0 84.8 89.7$\pm$2.6
M.S. MpMRI T2+ADC 20% patch 82.4 73.6 84.6 87.6$\pm$2.8
MpMRI T2+ADC 10% patient 75.3 80.7 73.9 86.1$\pm$2.9
M.S. MpMRI T2+ADC 10% patient 75.6 75.6 76.9 82.4$\pm$3.6
MpMRI T2+ADC 20% patient 72.5 74.3 72.0 81.5$\pm$3.6
M.S. MpMRI T2+ADC 20% patient 76.3 74.3 76.9 80.1$\pm$4.2
Table 3: Patch-based results of all performed experiments. Acc., Sen., Spec. and AUC are the shorts for Accuracy, Sensitivity, Specificity and Area Under (ROC)Curve, respectively. Standard deviations are computed from the 95% bootstrap confidence interval with $n=3000$ samples. The “Rate” column shows the filtering rate, and the “Method” column represents the filtering method. Best results for each section are in black bold; blue bold, and red bold separately. All units of the numeric values are in %.
Table 4 shows the original results based on image patches for the ablation study conducted in Section 6.3.
@@ A Original Results
Acc. Sen. Spec. AUC Acc. Sen. Spec. AUC
MpMRI 74.1 79.3 72.8 84.8$\pm$2.9 83.8 80.0 84.8 89.7$\pm$2.6
M.S. MpMRI 73.7 86.4 70.4 80.4$\pm$3.1 82.4 73.6 84.6 87.6$\pm$2.8
@@
Figure 7: Both figures demonstrate the original AUC results. 6(a) shows the comparison of AUC curves between the baseline and the models without filtration (experiments in the first category); 6(b) shows the comparison of AUC curves between the baseline, the best model without filtration, and the best model with filtration on the training set. “EFL” is short for Evidential Focal Loss. The shaded areas in both figures represent the 95% confidence intervals (CI) of each model. CIs are obtained by using Bootstrap with $n=3000$.
B Filtration while deploying
@@
((a))
((b))
Figure 8: Performance of filtering 0% to 40% of test data on the selected models. 7(a) represents the test performance of “MpMRI” model with 0% filtration on the training set, and 7(b) represents the test performance of “M.S. MpMRI” model with 0% filtration on the training set.

### v2

@@ 6.4 Filtering during deployment
So far, our exploration has centered on the impact of the data filtering strategy during training, aimed at enhancing model robustness and performance. It is also possible to apply filtering on the test set, i.e., when deploying the model to real clinical routines. This is equivalent to refraining from making decisions on the test samples that are identified as highly uncertain. The evaluation involves utilizing the pre-trained MpMRI and M.S. MpMRI models, each with 0% and 20% patch-based training filtering rates, as final models, and assessing their performance on the test set. The performance on filtering the test data based on different uncertainty thresholds using pre-trained models with 20% filtering during training is shown in Figure 4. For the performance of the other two models (0% filtering), refer to Appendix C. We started with an uncertainty threshold of 1.0 (keeping patients with ncertainty < 1.0); we progressively adjusted the threshold to 0.68, 0.63, 0.58, and ended at a threshold value of 0.53. These thresholds are determined based on the uncertainty value computed in the test set. We observe that the model improves its performance when filtering out highly uncertain patients from the test set, and eventually classified all patients correctly, as shown in Figure 4(a).
This pragmatic approach bears relevance in real clinical settings, offering radiologists an efficient means to allocate their time, focusing on patients filtered out during the diagnostic process due to their high uncertainty values, rather than those confidently classified.
((a)) MpMRI, with 20% filtering on training patches
((b)) M.S. MpMRI, with 20% filtering on training patches
Figure 4: Test performance of selected models based on uncertainty threshold.
@@ 7 Conclusion
In this study, we introduced a novel approach for unpaired image-to-image translation of prostate mp-MRI and developed a robust deep-learning model for classifying clinically significant PCa using evidential focal loss. We demonstrated the effectiveness of our method on our local dataset, reinforced by a publicly available one, and showed that uncertainty-aware filtering during both training and deployment can significantly improve the PCa classification performance. The quantitative results for image translation also demonstrated that our proposed domain transfer framework using the ACL-GAN model outperforms the baseline SDNet [5]. Our approach holds promise in expediting the diagnostic process by identifying patients with high uncertainty, allowing clinicians to focus on precise diagnosis and expedite cases with high prediction certainty.
While our approach has shown promising results, there are still opportunities for improvement. One potential area for future work is to consider the spatial dependency between slices in volumetric MRI. Currently, our domain transfer framework only accepts 2D images as input and output, and we reshape the volumetric MRI into several 2D slices. However, explicitly splitting 3D images into 2D slices may eliminate the spatial dependency within each MRI data and affect the classification results. Therefore, a plausible solution could be translating the entire 3D MRI volume from 3.0T to 1.5T instead of handling individual slices.
Lastly, there is great potential for further improving the classification performance by combining more images from different MRI functional sequences, such as b-value and Ktrans. We have already demonstrated that incorporating additional ADC images significantly enhances classification performance. We believe that if we successfully translate other images from b-value or Ktrans acquired at 3.0T to 1.5T and incorporate them into the classification, the results could be further improved. However, the additional MRI sequences may not be available in the local 1.5T dataset. The conversion process may become feasible if we acquire those sequences from local hospitals.
Disclosures
We declare we don’t have conflicts of interest.
Code, Data, and Materials Availability
ProstateX data is publicly available here  [^2] https://wiki.cancerimagingarchive.net/pages/viewpage.action?pageId=23691656, our local KHSC data is not available due to ethical concerns. Our code is at https://github.com/med-i-lab/DT_UE_PCa, we will make the code available upon acceptance.
@@ Acknowledgments
Appendix A Experimental Details
In this section, we provide details of hyperparameter settings to train uncertainty-aware classification models described in Section 5.1.
To train the “MpMRI” model for patch-driven filtering, we set the learning rate to 0.0001; weight decay to 0.01; total training epochs to 300, and batch size to 10. The class weights $\boldsymbol{\beta}$ in Equation (10) are set to [0.25, 0.75] for filtering 10%, and [0.25, 1.25] for filtering 20% of the training data. For patient-driven filtering, all parameters are the same except the class weights $\boldsymbol{\beta}$ are set to [0.25, 1] for filtering 10% of the training data. Last but not least, we set the initial learning rate to 0.0001; total training epochs to 300; batch size to 10; the learning rate decayed by a factor of 0.1 for every 200 epochs, and the class weights $\boldsymbol{\beta}$ are set to [0.25, 1] for filtering 20% of the training data.
Appendix B Original Results
In this section, we provide all original results based on image patches (Patch-based results) of all experiments we performed in the paper. We start by presenting Table 6, which is the patch-based result for Table 2 in Section 6.2.
Data Acc. Sen. Spec. AUC
SDNet [5](baseline) T2 77.8 27.9 90.7 74.8$\pm$4.1
T2-only T2 65.6 75.7 63.0 77.9$\pm$3.9
MpMRI T2+ADC 75.1 69.3 76.7 79.9$\pm$3.8
Vol. MpMRI T2+ADC 64.9 62.9 65.4 63.4$\pm$5.5
M.S. MpMRI T2+ADC 70.4 74.3 69.4 80.1$\pm$3.6
MpMRI+co-teaching T2+ADC 78.4 72.1 80.0 80.7$\pm$3.9
Table 6: Patch-based results of experiments using conventional training paradigm in Section 5.1. Standard deviations are computed from the 95% bootstrap confidence interval with $n=3000$ samples. The best results are bold. All units of the numeric values are in %.
Then, we provide the patch-based results for patch-based filtering in Table 7, which corresponds to Table 3 in Section 6.3.
Data F.R. F.M. Acc. Sen. Spec. AUC ECE $\downarrow$
MpMRI T2+ADC 10% patch 77.5 80.7 76.7 82.4$\pm$3.4 0.22
M.S.MpMRI T2+ADC 10% patch 79.7 70.7 82.0 81.1$\pm$3.8 0.16
MpMRI T2+ADC 20% patch 83.8 80.0 84.8 89.7$\pm$2.6 0.16
M.S. MpMRI T2+ADC 20% patch 82.4 73.6 84.6 87.6$\pm$2.8 0.17
Table 7: Patch-based results of experiments using evidential focal loss and patch-based filtering. Standard deviations are computed from the 95% bootstrap confidence interval with $n=3000$ samples. The “F.R.” and “F.M.” represent the filtering rate and the filtering method, respectively. The best results for each filtration rate are bold. All units of the numeric values are in %.
For patient-based filtering, we provide the patch-based results in Table 8, which corresponds to Table 4 in Section 6.3.
Data F.R. F.M. Acc. Sen. Spec. AUC ECE $\downarrow$
MpMRI T2+ADC 10% patient 75.3 80.7 73.9 86.1$\pm$2.9 0.21
M.S. MpMRI T2+ADC 10% patient 75.6 75.6 76.9 82.4$\pm$3.6 0.17
MpMRI T2+ADC 20% patient 72.5 74.3 72.0 81.5$\pm$3.6 0.22
M.S. MpMRI T2+ADC 20% patient 76.3 74.3 76.9 80.1$\pm$4.2 0.18
Table 8: Patch-based results of experiments using evidential focal loss and patient-based filtering. Standard deviations are computed from the 95% bootstrap confidence interval with $n=3000$ samples. The “F.R.” and “F.M.” represent the filtering rate and the filtering method, respectively. The best results for each filtration rate are in bold. All units of the numeric values are in %.
Table 9 shows the original patch-based results based on image patches for the ablation study conducted in Section 6.3.
@@ Appendix B Original Results
Acc. Sen. Spec. AUC ECE Acc. Sen. Spec. AUC ECE
MpMRI 74.1 79.3 72.8 84.8$\pm$2.9 0.21 83.8 80.0 84.8 89.7$\pm$2.6 0.16
M.S. MpMRI 73.7 86.4 70.4 80.4$\pm$3.1 0.22 82.4 73.6 84.6 87.6$\pm$2.8 0.17
@@
Figure 5: Both figures demonstrate the original AUC results. 5(a) shows the comparison of AUC curves between the baseline and the models without filtration (experiments in the first category); 5(b) shows the comparison of AUC curves between the baseline, the best model without filtration, and the best model with filtration on the training set. “EFL” is short for Evidential Focal Loss. The shaded areas in both figures represent the 95% confidence intervals (CI) of each model. CIs are obtained by using Bootstrap with $n=3000$.
Appendix C Filtration while deploying
@@
((a)) MpMRI, with 0% filtering on training patches
((b)) M.S. MpMRI, with 0% filtering on training patches
Figure 6: Test performance of selected models based on uncertainty threshold.
First Author insert biography
Biographies and photographs of the other authors are not available.